What is the theoretical yield (in grams) of C9H8O4 when 2.00 g C7H6O3 is heated
with 4.00 g of C4H6O3? If the actual yield is 1.98 g, what is the percent yield?
C7H6O3 + C4H6O3 → C9H8O4 + C2H4O2
1
Expert's answer
2012-09-13T12:08:43-0400
C7H6O3 + C4H6O3 → C9H8O4 + C2H4O2

n of C7H6O3 = 2/ (12*7 + 6 + 3*16) = 0.0145 moles this compound is less

n of C4H6O3 = 4/(12*4 + 6+3*16) = 0.0392 moles

0.0145 0.0145
C7H6O3 + C4H6O3 → C9H8O4 + C2H4O2

1. m of C9H8O4 = n* Mw = 0.0145* (12*9 + 8+ 16*4) = 2.61 g

2. w = 1.98/2.61*100% = 75.86%